The perimeter of a rectangle wildlife perserve is 8 kilometers, and its area is 3 square kilometers. Find its dimensions.
Answer by rajagopalan(174)

let one side =a
let other side =b
perimeter=2(a+b)
2(a+b)=8(given)
a+b=8/2=4
so b=4-a
area = a*b
area=a*(4-a)
area=4-a^2
this is equal to 3 sq km
4-a^2=3
4-3=a^2
1=a^2
a=1
b=4-a
b=4-1=3
the dimns are 3km and 1 km
good day